Circular No. 7709

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V4740 SAGITTARII F. Patat, European Southern Observatory; C. Contreras and J. Prieto,

Pontificia Universidad Catolica de Chile; and G. Altavilla, S. Benetti, E. Cappellaro, A. Pastorello, and M. Turatto, Osservatorio Astronomico di Padova, report: “A low-resolution spectrum (range 320–900 nm, resolution 1.2 nm) of this object (cf. IAUC 7706, 7708) was obtained on Sept. 9.238 UT using the Danish 1.5-m telescope (+ DFOSC) at La Silla. The spectrum is dominated by strong emission lines of the Balmer series of H I and by lines of Fe II of multiplets 42, 49, and 74. The FWHM measured from the Fe emissions is ∼ 1200 km/s. The Na I D line has a P-Cyg profile with the minimum of the absorption corresponding to 1500 km/s. In general, the spectrum is very similar to that of the 1989 nova V443 Sct soon after maximum (Rosino et al. 1989, A.J. 101, 1807).”

Photometry by A. C. Gilmore, Mount John Observatory, University of Canterbury (reference stars Cousins E745 and E746): Sept. 8.416 UT, V = 6.71, U − B = +0.31, B − V = +0.82, V − R = +0.45, V − I = +0.88, air mass 1.15; 9.333, 7.36, −0.01, +0.69, +0.47, +1.07, 1.03. Broadband V magnitudes by W. Liller, Viña del Mar, Chile: Sept. 7.996, 7.04; 8.987, 7.44. Visual magnitude estimates: Sept. 8.781, 7.5 (S. Baroni, Milan, Italy); 8.812, 6.9 (M. Reszelski, Szamotuly, Poland); 9.042, 7.1 (J. G. de S. Aguiar, Campinas, Brazil); 9.481, 7.5 (A. Pearce, Nedlands, W. Australia).

NOVA IN M31 M. Fiaschi, F. Di Mille, and R. Cariolato, Astronomical Observatory

G. Colombo, Padua, report their discovery of an apparent nova at α = 0^{h}42^{m}34^{s}.55, δ = +41^{o}18′13′′.8 (equinox 2000.0), which is ∼ 146′′ west and 124′′ north of the center of M31. The CCD images were obtained with an Hα filter (6.3-nm FWHM) and a 0.41-m reflector, yielding the following magnitudes: Sept. 2.93 UT, 15.9; 3.90, 15.8. Nothing was visible at this position on an image taken on Aug. 27.9 (limiting mag 17.5).

SUPERNOVA 2001ed IN NGC 706 W. M. Wood-Vasey, University of California at Berkeley and Lawrence

Berkeley National Laboratory, reports pre-discovery observations of SN 2001ed (cf. IAUC 7703, 7704) taken with the 1.2-m NEAT/MSSS telescope at Haleakala on Aug. 25.32 and 25.36 UT, which clearly show the new object in a knot in the spiral arm of NGC 706 at unfiltered mag 15.85.

2001 September 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green